The total electrical flow through a closed surface similar to a cylinder is 1.1 x 104 N m2 / C. The net load on the cylinder is:
a.18 µC
b.45 nC
c.72 nC
d.97 nC
Answer #1

here,

The total electrical flow through a closed surface similar to a cylinder , phi = 1.1 * 10^4 N m2 / C.

let the net load in the cyclinder be q

the net electric flux , phi = q /e0

1.1 * 10^4 = q / (8.85 * 10^-12)

q = 97 * 10^-9 C = 97 nC

the net load on the cyclinder is d) 97 nC